OpenStudy (anonymous):

how do I find the equation with the following zeros: -1, 4-5i

OpenStudy (anonymous):

multiply out (expand) the left side if you want standard form.

OpenStudy (anonymous):

Where did the 3 come from? this is the answer I came up with x^3-7x^2+33x+41 for the first part that you show I came up with (x+1)(x-4+5i)(x-4-5i). Is this correct?

OpenStudy (anonymous):

I have no idea!!! it's supposed to be the -1.

OpenStudy (anonymous):

since you list a complex root, it's conjugate is also a root so you actually have 3 roots: x = -1, x = 4-5i, x = 4+5i (conjugate)

OpenStudy (anonymous):

lol I was freaking out over here!

OpenStudy (anonymous):

so now just create your equation with the listed as your solution: (x--1)*(x-(4-5i))(x+(4+5i)) = 0

OpenStudy (anonymous):

sorry, bout that!

OpenStudy (anonymous):

no problem, I don't fully grasp all of the steps, but I think I'm doing most of it correctly, I just don't know where I am going wrong.

OpenStudy (anonymous):

if you know the zeros to a function, always write it out in factored form then expand it.

OpenStudy (anonymous):

lost internet connection for a while....

OpenStudy (anonymous):

so you should get: x^3 - 7x^2 + 33x + 41 = 0

OpenStudy (anonymous):

my screen froze

OpenStudy (anonymous):

ok, I stumbled through it but that's what I came up with, it's just confusing.

OpenStudy (anonymous):

do I need to put the =0 part?